Solve log{10}(.04) when log{10}(5)=.7
Found 2 solutions by Edwin McCravy, stanbon:
Answer by Edwin McCravy(20059)
Answer by stanbon(75887)

Solve log{10}(.04) when log{10}(5)=.7
-------
0.04 = 4/100 = 1/25 = 1/5^2 = 5^-2
---------------------------------------------
Therefore log(0.04) = log(5^(-2))
= -2log(5)
=-2*0.7
= -1.4
===========================
Cheers,
Stan H.